OpenStudy (anonymous):

In a single throw of a die, what is the probability of attaining a number of 4 or lower?

OpenStudy (kainui):

How many possibilities are there? 6 How likely are you to get any number? They're all equally likely. So each number has a one out of six chance of getting picked. We write that as 1/6 for any one you pick. So what's the probability of rolling either 1 or 2? Well, you have a 1/6 chance of getting a 1 and a 1/6 chance of getting a 2. So you add them up and get \[\frac{1}{6}+\frac{1}{6}=\frac{2}{6}=\frac{1}{3}\] Now back to your question, what is the probability of 4 or lower? Well, just add up the individual probabilities of getting each one.

OpenStudy (anonymous):

How can you get 4 or less? Well, you can get 1 or 2 or 3 or 4. Anytime you can answer the probability qestion with “or” you add the individual probabilities: P(4 or less) = P(1) + P(2) + P(3) + P(4) Each of these events is equally likely, so P(4 or less) = 1/6 + 1/6 + 1/6 + 1/6 = 4/6 = 2/3

OpenStudy (anonymous):

0.67

OpenStudy (anonymous):

?

OpenStudy (anonymous):

2/3 is approximately .67